Round to 4th decimal place, if needed. Use your calculator to find the 5-number summary of the dataset: 143,15,7,47,3,48,40,18,45,49) What is the Q1? 2. What is the Median? 3.What is the Q3? 4. What is the IQR?
Answer #1

Solution:

Given that,

Arranging Observations in the ascending order,
3,7,15,18,40,43,45,47,48,49

n = 10

1 ) Q1

Q1 = ( n + 1 / 4 )th value of the observation

= ( 10 + 1 / 4 )th value of the observation

= ( 11 / 4 )th value of the observation

= 2.75th value of the observation

= 2 nd Observation +0.75 ( 3rd + 2 nd )

= 7 + 0.75 ( 16 - 7 )

= 7 + 0.75( 8 )

= 7 + 6

= 13

Q1 = 13

2 ) Median

M = ( n / 2 )th value of the observation + ( n /2 + 1 )th value of the observation ) / 2

= (10 / 2 )th value of the observation + ( 10 /2 + 1 )th value of the observation ) / 2

= (5 )th value of the observation + ( 6 )th value of the observation ) / 2

= 2.75th value of the observation

= 40 + 43 / 2

= 83 /2

= 41 .5

Median = 41.5

1 ) Q3

Q3= ( 3*n + 1 / 4 )th value of the observation

= ( 3 * 11 / 4 )th value of the observation

= ( 33 / 4 )th value of the observation

= 8.25th value of the observation

= 8th Observation +0.25 ( 9th + 8th )

= 47 + 0.25 ( 48 - 47 )

= 47 + 0.25( 1 )

= 47 + 0.25

= 47.25

Q3 = 47.25

4 ) Inter Quartile range is IQR

IQR = Q3 - Q1

= Upper Quartile - Lower Quartile

= 47.25 - 13

= 34.25

IQR = 34.25